OpenStudy (anonymous):

Help needed here...? Which of the following tables of values satisfies the equation below? A. t b 2 20 4 40 6 60 B. t b 2 12 4 22 6 32 C. t b 2 12 4 14 6 16 D. t b 2 14 4 16 6 18

OpenStudy (anonymous):

Sorry, the equation is b = 10 + t

OpenStudy (katrinakaif):

Lets plug the values given in 'A' in, and see what you get

OpenStudy (anonymous):

So the numbers in A, under "b" are the only ones that is divisible by 10?

OpenStudy (katrinakaif):

You have to plug both values for b and t in. In the first case, b = 20 while t = 2 Plug in: b = 10 + t 20 = 10 + 2 Is it true?

OpenStudy (anonymous):

No.. it is 20 = 10 x 2

OpenStudy (anonymous):

My next choice would be C.

OpenStudy (katrinakaif):

Your equation is : b= 10 + t or b = 10 x t?

OpenStudy (anonymous):

b = 10 + t

OpenStudy (katrinakaif):

Yes so that makes it invalid, now Choice C, Lets plug it in.

OpenStudy (katrinakaif):

Now you have to check for all values.

OpenStudy (anonymous):

b = 10 + t 12 = 10 + t 12 = 10 + 2 ✔

OpenStudy (anonymous):

b = 10 + t 14 = 10 + t 14 = 10 + 4 ✔ b = 10 + t 14 = 10 + t 14 = 10 + 6 ✔ My final answer is C and I am confident in my answer. ✔

OpenStudy (katrinakaif):

Great! :-)

OpenStudy (anonymous):

Thank you so much for guiding me step by step! :D

OpenStudy (katrinakaif):

The pleasure is mine :-)